If the speed of sound depended on frequency, what problems would this create for live concerts?

A light string is attatched to a heavy rope, and the whole thing is pulled tight.   A wave is sent along the light string.   When it hits the heavy rope,  compared to the wave on the string, the wave that propagates along the rope has the same (A) frequency (B) wavelength (C) both frequency and wavelength (D) neither?
Answer #1

we have ralation

Velocity \(V=f \lambda\)

frequency for both the medium ramian same so

\(\frac{V_{w}}{\lambda_{w}}=\frac{V_{a}}{\lambda_{a}}\)

here \(V_{w}\) is speed of sound in water \(, V_{a}\) is speed of sound in air. \(\lambda_{w}\) is wavelength in water and \(\lambda_{a}\) is wavelength in air.

putting all values

\(\frac{V_{w}}{5.76 \times 10^{-3}}=\frac{343}{1.36 \times 10^{-3}}\)

\(\mathrm{V}_{\mathrm{W}}=1452.70 \mathrm{~m} / \mathrm{s} \)

answered by: ---